20181029 UT

Proposed Schedule (seeing <1.2" on DIMM - transparency up to thin clouds):

  1. DONE but clouded out - LBTI - LMIRCAM/NOMIC TOO observations of Comet (3 hours)
  2. DONE On-sky checkout of AGW-1 (1-3 hours)
    • confirm light on lenslets
    • measure/update the probe wrt/rotator center
  3. DONE but some discrepancies - LUCI-1 measure rotator zeropoint - old method
  4. DONE LUCI-1 AGW Transform
    • Transform field candidates:
      • Stone O sunset until 10pm (30 deg limit at 11pm)
      • Stone A1 sunset until 2am (30 deg limit at 3am)
      • Stone D 11pm onwards
  5. Confirm LUCI-1 AGW Transform
  6. MODS Update rotator zero-point for MODS-2 (MODS-1 should be OK)
  7. MODS AGW Transform - requies pointing mask (1 hour)
  8. Time Permitting - D-time Science (MODS-LS Triple AGN Candidate - Rothberg, and/or NOVA stars - Wagner)

Poor Weather Backup:

  1. collect pointing data with MODS/MODS (~2.5 hours, any seeing)
  2. DONE TMS DX M1 Tilt test (10 min)
  3. DX M3 alignment at other focal station
    • for PFU (high)
    • for L-N (high)
  4. Test binocular and pseudomonocular LBC+LUCI/MODS observations:
    • Can we send binocular LBC-Red + LUCI-1 presets via scripts (Olga, a bad-seeing, bad-weather task)
    • Can we send binocular LBC-Red + MODS-1 binocular presets via scripts (Olga, a bad-seeing, bad-weather task)
    • As far as pseudomonocular mode, can IT 5715 be closed? (we believe both are fixed in version 2.15, which is the current version of the LBC software. when dec=-90, is an initial, non-zero, dither made, or is it ignored? do LBC observations always wait for the optics move to complete?
    • DONE Test new version of GCS
  5. H-Band Library (LUCI-Bino) - Poor Weather Science
  6. Radio Quiet AGN Sample (MODS or LUCI) - Poor Weather Science

Details

UT 01:26 - Aux Cryo2 has successfully been removed from telescope. LUCI-1 has no MOS masks in cabinet.

UT 01:27 - experiencing some HBS issues with Inverter for Pump #4

UT 01:45 - LUCI-1 MOS xml updated to reflect no MOS masks in LUCI-1.

UT 01:57 - HBS onine - doing LBC fills

LBTI 29P

UT 02:31 - Open and handed over to LBTI

UT 02:44 - DIMM shows 1.73"

UT 02:48 - running init all on LUCI-1

UT 02:54 - inti all on both LUCIs ok, 0001-0002 L1/L2 2.51 sec dark images

UT 05:22 - LBTI giving up - configuring for LUCI-Bino

LUCI/LUCI for AGw checkout

UT 05:50 Adjusting pointing on bright star - clouds starting to clear a bit. SX IE=-64 CA=+44 DX IE=-20 CA=15

UT 06:xx ACTIVE optics converges from a nasty starting point. PSF Left was set to "secondary" rather than "split" causing GCS to complain that PSF rejected Zernikes (which is somewhat correct, but not very explanatory). Worked fine after that. Seeing is variable in the 0.7-1 arcsec range.


LUCI/LUCI probe centering

UT 06:17 SX Stop Guiding; Homed probe; setxy 0,612.5; +5 CA

/lbt/data/share/tcs/lbto.20181029/left_guider000002.fits

UT 06:22 SX circle is 0002 measured 116,112, unbinned 232,224, detsec 247,263 (not very round) hotspot 242,256 = close enough

UT 06:25 DX Stop Guiding; Homed probe; setxy 0,612.5

/lbt/data/share/tcs/lbto.20181029/right_guider000002.fits

UT 06:29 DX circle is 0002 measured 131,114, unbinned 262,228, detsec 275,245 (nicely round) hotspot 275,243 = OK

UT 06:35 resend ACTIVE preset.

UT 06:40 ACQUIRE preset for pointing check because we lost the DX star. Was that because I homed the probe? Hmmmmm.

SX IE=-54 CA=+39 DX IE=-25 CA=+15 values didn't change so much from earlier?

LUCI/LUCI rotator zeropoint

UT 07:02 ACTIVE preset to STONEA1 gs=2 R=14.7 (gs=0 was too faint at R=17.0)

UT 07:08 left_wfscimage000069.fits when stretched looks like it has poor coma correction, but the guider image was nice ???

UT 07:12 TRACK preset for offsets along slit

UT 07:18 - completed, but stars might be too bright, taking test exposure with [FeII] filter luci1.20181029.0014/luci2.

UT 07:20 - running with [FeII] version of script - luci.1.20181029.0015-0025,

UT 07:29 - putting mask in FPU - luci1/2.20181029.0025 thru-slit (0.5")

UT 07:30 - LUCI-2 MOS error - Mask 8 cannot make it into FPU - taking thruslit for LUCI-1, then continuing, will also diagnose MOS error

UT 07:34 - done with data - John Hill will work on other sub-systems

UT 07:40 - deployed patches for PCS, PSF and GCS

UT 08:00 - changed LUCI LEFTZEROPOINT from -187.917 to -188.519 and restarted PCS again to read PCSInstrument.conf. - based on analysis from slitscan_luci using 11 data points (need to add -0.6 to rotator)

UT 08:11 - re-authorizing to LUCI-1 - sending mono version of script - verifying the offset added to PCS

luci1.20181029.0027-0037-0030 - image quality went bad - stars aren't round, paused script - Josh sending active preset to collimate then we'll redo script

UT 08:20 - sending track script again

luci120181029.0031-0041

UT 08:25 - Jenny has diagnosed the LUCI-2 MOS error problem - notes to follow - LUCI-2 MOS recovered

UT 08:45 - new resutls after putting in -0.6 rotator is that we need -0.43 added to the rotator center

UT 08:48 - moving to Stone D which is at higher elevation to repeat this test.

UT 08:51 - staying monocular because clouds are moving in, trying to squeeze in this L1 measurement

luci1.20181029.0042-0052

UT 09:12 - new field (Stone D) shows offset of -1.67 is needed, which doesn't make sense, going back to Stone A data, 2nd star in that field shows -0.59 (consistent)

UT 09:17 - looks like tilt is more to the right on Stone D field so this suggests we got the sign wrong, changing this to +0.6 offset in PCS and redoing it

UT 09:20 - changed LUCI LEFTZEROPOINT to -187.315 deg , changed filter to H-band for script (remaining monocular for now)

UT 09:22 - taking thru-slit LUCI-2 since I already took data - wish me luck!
UT 09:25 - luci2.20181029.0027 - thru slit

UT 09:27 - running LUCI-1 monocular confirmation check while I analyze LUCI-2 data

luci1.20181029.0053-0063

UT 09:31 - LUCI-2 DX shows -0.14 rotation ZP offset (should this be +0.14 in PCS?) - not updating for now - will leave as is

UT 09:39 - change to LUCI-1 looks a lot better - the analysis claims +0.49 needs to be added but the vertical alignment of slit and stars look very close by eye - will run Transform now.

LUCI-1/-2 AGW Transform Collection

UT 09:40 - authorize binocular LUCI

UT 09:41 - pointing check first

UT 09:48 putting N30 Fieldstop in both LUCIs

Setting up for N3.75 Camera, FeII filter, 2.51 sec

-need to take a dark

luci1.20181029.0064 - 2.51sec x2 NDIT Dark, 0065-0066 - checking fieldstop alignment

luci2.20181029.0028 - 2.51sec x2 NDIT Dark, 0029 - checking fieldstop alignment

N30 fieldstops look OK

UT 09:59 - Ready to go on the LUCI-side, waiting for pointing

UT 10:12 Start: DIMM 0.72"

L2-0030

UT 10:14 - AGW1 conked out with:

Errroor in station check: AGWNAME from header not found in station_params file: left_wfscimage 000132.fits

-could be AGW4 still hard-coded somewhere? John Hill checking

John found the error - updaed an IDL script

UT 10:21 - started AGW1 transform

-wfscimages not showing anything

UT 10:30 - may not be collimating on SX, but seeing is so good, its not affecting things - John is hunting down issue while we continue to take data

UT 10:31 - John found problem, but continuing - issue is wrong area is cut out for SH analysis (flip from AGW4) so its not collimating and not finding spots

UT 10:37 - stopped SX at L1-0075 IGNORE ALL LUCI-1 DATA FROM 0067-0075

UT 10:38 - restarted LUCI-1 AGW-1 SX - DIMM 0.79"

L1-0076

UT 10:40 - lost GS on DX side - stopped at L2-0044 - Josh nudging star back in (GS23 last star completed)

Restarting DX #16, (start=16)

UT 11:10 - DX failed with GS30 PA=30 failed to move position - out of range 178.68, -1.47361mm (x,y)

-restarting at next number - last one in the list

Finished DX Side

UT 11:24 - SX lost guidestar GS12 PA=30

UT 11:25 - resarting SX at GS12 PS=30

UT 11:27 - Failed at GS 23 PA=30, skipping to GS 13 PA=30

UT 11:32 - failed at GS5 PA=30, skipping to GS 11 PA=30

UT 11:36 - failed at GS29 - stopping

L1 luci1.20181029.0076-0104, darkframe is luci1.20181029.0064

L2 luci2.20181029.0030-0061, darkframe is luci2.20181029.0028

STOPPING LUCI-Bino - putting LUCIs in safe mode

Transform Analsys 29 Oct 2018 Doug

DX

  • LBTO@obs2:data/20181029/rfbg/transform_rfbg.log
  • LUCI2 image
    • Start star 4 @ 10:16 /Repository/20181029/luci2.20181029.0032.fits
    • End star 24 @ 11:09 /Repository/20181029/luci2.20181029.0060.fits
  • ~/work/transform/rfbg/20181029
    • rfbg.log
    • probe_position.dat (probe positions from TCS log. tramsform_luci2.log values for probe are not correct)
  • IDL> transform_process_set, 'rfbg.log', set, transform, retval, /fwhm, /pt_mask, /display

SX

  • LBTO@obs2:data/20181029/lfbg/transform_lfbg.log
  • LUCI1 image
    • Start star 4 @ 10:40 /Repository/20181029/luci1.20181029.0076.fits
    • End star 11 @ 11:36 /Repository/20181029/luci2.20181029.0104.fits
  • ~/work/transform/lfbg/20181029
    • lfbg.log
    • probe_position.dat (probe positions from TCS log. tramsform_luci1.log values for probe are not correct)
  • IDL> transform_process_set, 'lfbg.log', set, transform, retval, /fwhm, /pt_mask, /display


UT 11:38 - Reconfiguring for MODS

UT 11:44 - waking MODS-1 MODS-2

UT 11:48 - Red Grating on MODS-1 shows error - red box in GUI and Sievesnap produced an odd exposure, log showed

M1.IE > MC1 ERROR RGRATING = TIMEOUT cannot read from 192.168.139.101:8003

Red Dewar pressure was also N/A - restarted IC but pressures are still not reporting

Following troubleshooting on:

https://wiki.lbto.org/Instrumentation/TroubleshootingISS to connect to IP of 192.168.139.101 to reboot

After rebooting the comtrol ports, it is necessary to restart the IE service (mods1 start ie) on mods1data

-IT 6877 - escalated to Critical

TMS M1 displacement

UT 12:10 Did Andrew's DX M1 displacement test to cross-check the Gregorian TMS results.

"The test is to get a bright star; note its pointing, de-centre the M1 by 1 mm, and look to see if ~ 8 arc seconds of pointing change has occurred at the coma tip (or what if any change anyway..."

Result: I did this test just now at RDG. 0.1 mm of DX M1 decenter (X or Y Global Offset of M1) displaces the image by ~3 arcsec. Same result for X and Y decenters with expected orthogonality. It seems repeatable within telescope drift. That was more motion than Andrew was expecting.

Close

UT 12:17 - put MODS to sleep while John was working on AGWs......unilateral decison I guess....

UT 12:20 Close for impending dawn, and sky starts to clear within minutes.

-- BarryRothberg - 26 Oct 2018
Topic revision: r10 - 07 Jun 2019, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback